OpenStudy (anonymous):

A car travels 1/6 of the distance between two cities in 3/5 of an hour. At this rate, what fraction of the distance between the two cities can the car travel in 1 hour?

OpenStudy (anonymous):

I think its 5/15

OpenStudy (taramgrant0543664):

It's not that the top number is right but the bottom is not

OpenStudy (anonymous):

Ok can you guide me threw the steps?

OpenStudy (taramgrant0543664):

(1/6) / (3/5)=x/1 Unfortunately I'm on my iPad so I can't use the equation thing but this is the initial set up

OpenStudy (anonymous):

5/18?

OpenStudy (taramgrant0543664):

When you have a number divided by a number that is divided you flip the bottom and multiply (1/6) x (5/3)=x

OpenStudy (taramgrant0543664):

And yep you'll get 5/18!

OpenStudy (anonymous):

thank you:)